# Evolution of cooperation in a bimodal mixture of conditional cooperators

### v1

@@ Abstract
Extensive behavioral experiments reveal that conditional cooperation is a prevalent phenomenon. Previous game-theoretical studies have predominantly relied on hard-manner models, where cooperation is triggered only upon reaching a specific threshold. However, this approach contrasts with the observed flexibility of human behavior, where individuals adapt their strategies dynamically based on their surroundings. To capture this adaptability, we introduce a soft form of conditional cooperation by integrating the Q-learning algorithm from reinforcement learning. In this framework, players not only reciprocate mutual cooperation but may also defect in highly cooperative environments or cooperate in less cooperative settings to maximize rewards. To explore the effects of hard and soft conditional cooperators, we examine their interactions in two scenarios: structural mixture (SM) and probabilistic mixture (PM), where the two behavioral modes are fixed and probabilistically adopted, respectively. In SM, hard conditional cooperators enhance cooperation when the threshold is low but hinder it otherwise. Surprisingly, in PM, the cooperation prevalence exhibits two first-order phase transitions as the probability is varied, leading to high, low, and vanishing levels of cooperation. Analysis of Q-tables offers insights into the “psychological shifts” of soft conditional cooperators and the overall evolutionary dynamics. Model extensions confirm the robustness of our findings. These results highlight the novel complexities arising from the diversity of conditional cooperators.
@@ I Introduction
Recently, reinforcement learning (RL) has emerged as a new paradigm Kaelbling et al. (1996); Rangel et al. (2008); Lee et al. (2012) for understanding human behavior, offering a flexible approach to modeling conditional cooperation. Unlike hard conditional cooperators (HCC), RL-based players make decisions guided by policies that are continuously revised to maximize accumulated rewards. In this way, each player develops a unique policy in an introspective manner. To date, several studies have combined RL with evolutionary game theory to investigate various aspects of human behavior Tomov et al. (2021); Zhang et al. (2020a); Wang et al. (2022); Sheng et al. (2024); Jia et al. (2021); Ding et al. (2024), including cooperation Zhang et al. (2020b); Song et al. (2022); Wang et al. (2023); Ding et al. (2023); Zhao et al. (2024a); Ma et al. (2023); Zheng et al. (2024a); Li et al. (2025), trust Zheng et al. (2023a), fairness Zheng et al. (2024b), resource allocation Zhang et al. (2019); Zheng et al. (2023b); He et al. (2022), and other collective behaviors Zhang et al. (2020c); Ding et al. (2024). Notably, Refs. Ezaki et al. (2016); Horita et al. (2017) demonstrate that RL explains conditional cooperation and moody conditional cooperation. More recent studies Geng et al. (2022); Zhao et al. (2024b); Sheng et al. (2024) have revealed that RL players exhibit features of conditional cooperation, cooperating in cooperative environments and defecting in defective ones. Interestingly, they also display hump-shaped conditional cooperation, where cooperators switch to defection in highly cooperative neighborhoods to maximize rewards. Given the ubiquity of conditional cooperation and the potential of RL, it is intriguing to incorporate this new RL perspective to describe conditional cooperative behaviors beyond traditional models. The central question we address is: How does cooperation evolve when diverse conditional cooperation modes are incorporated?
@@ II Model
Figure 2: Evolution of cooperation with structure mixing. (a) Time series of cooperation level $f_{C}$ for different proportion $\rho$, where the threshold for HCC $\xi=1,3,4$ for the three subplots, respectively. (b) The prevalence of cooperation $f_{c}$ as a function of $\rho$ for $\xi=0,1,...,4$. (c) Phase diagram for $f_{C}$ in the parameter domain of $\rho-b$ with $\xi=3$. Other parameter: $b=0.1$ in (a,b).
@@
This means that player $i$ chooses the suggested action $\underset{a\in A}{\arg\max Q(s_{i}(t),a)}$ with a probability of $1-\epsilon$, and makes a random exploration with $\epsilon$, where $0<\epsilon\ll 1$.
Figure 3: Evolution of cooperation with probabilistic mixing. (a) Time series of cooperation level $f_{C}$ for different $\rho$, where the threshold $\xi=3$. The insert shows the probability density function (PDF) distribution of $f_{C}$ within the steady states. (b) $f_{C}$ as a function of the probability $\rho$ for $\xi=0,1,...,4$. The two insets show the zoom-in of the two transition regions. (c) Phase diagram for $f_{C}$ in the parameter domain $\rho-b$ with $\xi=3$. Other parameter: $b=0.1$ in (a,b).
@@ IV.2 Probabilistic mixing
Figure 8: Evolution of Q-tables in PM. Left column: The temporal evolution of averaged Q-value difference $\overline{\Delta Q}_{s_{j}}$ defined in Eq. (8) for all six states. Right column: The temporal evolution of all six state densities $f_{s_{0,...,5}}$. Parameter for the three rows: (a, d) $\rho=0.3$, (b, e) $\rho=0.4$, (c, f) $\rho=0.5$. Other parameter: $\xi=3$.
@@
$\bar{f}\left(s,s^{\prime}\right):=\frac{\sum_{\tau=t_{0}}^{t-1}\sum_{i\in\mathcal{N}}\mathbbm{1}_{s^{i,i}(\tau)=s,s^{i,i}(\tau+1)=s^{\prime}}}{|\mathcal{N}|\left(t-t_{0}-1\right)}.$ (10)
Therefore, Cohen’s kappa coefficient $\kappa(s,s^{\prime})$ captures the transition “flow” from state $s$ to $s^{\prime}$. If a pair of non-diagonal positive coefficients $\kappa(s,s^{\prime})\approx\kappa(s^{\prime},s)$, this corresponds to a detailed balance, otherwise there is a net transition flow between the two states. Note that diagonal elements of the matrix are for self-loops, where the system could settle down.
@@ VI Conclusion and discussion
In both mixing schemes, Q-learning players act as a “cooperation lubricant” in scenarios where $\xi=3$, where cooperation would otherwise fade due to the inability of hard conditional cooperators alone to sustain it. Remarkably, through Q-learning, players cooperate to achieve higher long-term rewards, ultimately fostering a significant level of cooperation in environments that would typically favor defection.

### v2

@@ Abstract
Extensive behavioral experiments reveal that conditional cooperation is a prevalent phenomenon. Previous game-theoretical studies have predominantly relied on hard-manner models, where cooperation is triggered only upon reaching a specific threshold. However, this approach contrasts with the observed flexibility of human behaviors, where individuals adapt their strategies dynamically based on their surroundings. To capture this adaptability, we introduce a soft form of conditional cooperation by integrating the Q-learning algorithm from reinforcement learning. In this form, players not only reciprocate mutual cooperation but may also defect in highly cooperative environments or cooperate in less cooperative settings to maximize rewards. To explore the effects of hard and soft conditional cooperators, we examine their interactions in two scenarios: structural mixture (SM) and probabilistic mixture (PM), where the two behavioral modes are fixed and probabilistically adopted, respectively. In SM, hard conditional cooperators enhance cooperation when the threshold is low but hinder it otherwise. Surprisingly, in PM, the cooperation prevalence exhibits two first-order phase transitions as the probability is varied, leading to high, low, and vanishing levels of cooperation. Analysis of Q-tables offers insights into the “psychological shifts” of soft conditional cooperators and the overall evolutionary dynamics. Model extensions confirm the robustness of our findings. These results highlight the novel complexities arising from the diversity of conditional cooperators.
@@ I Introduction
Recently, reinforcement learning (RL) has emerged as a new paradigm Kaelbling et al. (1996); Rangel et al. (2008); Lee et al. (2012); Wang et al. (2024) for understanding human behavior, offering a flexible approach to modeling conditional cooperation. Unlike hard conditional cooperators (HCC), RL-based players make decisions guided by policies that are continuously revised to maximize accumulated rewards. In this way, each player develops a unique policy in an introspective manner. To date, several studies have combined RL with evolutionary game theory to investigate various aspects of human behavior Tomov et al. (2021); Zhang et al. (2020a); Wang et al. (2022); Sheng et al. (2024); Jia et al. (2021); Ding et al. (2024), including cooperation Zhang et al. (2020b); Song et al. (2022); Wang et al. (2023); Ding et al. (2023); Zhao et al. (2024a); Ma et al. (2023); Zheng et al. (2024a); Li et al. (2025); Mintz and Fu (2025), trust Zheng et al. (2023a), fairness Zheng et al. (2024b), resource allocation Zhang et al. (2019); Zheng et al. (2023b); He et al. (2022), and other collective behaviors Zhang et al. (2020c); Ding et al. (2024). Notably, Refs. Ezaki et al. (2016); Horita et al. (2017) demonstrate that RL explains conditional cooperation and moody conditional cooperation. More recent studies Geng et al. (2022); Zhao et al. (2024b); Sheng et al. (2024) have revealed that RL players exhibit features of conditional cooperation, cooperating in cooperative environments and defecting in defective ones. Interestingly, they also display hump-shaped conditional cooperation, where cooperators switch to defection in highly cooperative neighborhoods to maximize rewards. Given the ubiquity of conditional cooperation and the potential of RL, it is intriguing to incorporate this new RL perspective to describe conditional cooperative behaviors beyond traditional models. The central question we address is: How does cooperation evolve when diverse conditional cooperation modes are incorporated?
@@ II Model
Figure 2: Evolution of cooperation with structure mixing (SM). (a) Time series of cooperation level $f_{C}$ for different proportion $\rho$, where the threshold for HCC $\xi=1,3,4$ for the three subplots, respectively. (b) The prevalence of cooperation $f_{c}$ as a function of $\rho$ for $\xi=0,1,...,4$. (c) Phase diagram for $f_{C}$ in the parameter domain of $\rho-b$ with $\xi=3$. Other parameter: $b=0.1$ in (a,b).
@@
This means that player $i$ chooses the suggested action $\arg\underset{a\in A}{\max}Q(s_{i}(t),a)$ with a probability of $1-\epsilon$, and makes a random exploration with $\epsilon$, where $0<\epsilon\ll 1$.
Figure 3: Evolution of cooperation with probabilistic mixing (PM). (a) Time series of cooperation level $f_{C}$ for different $\rho$, where the threshold $\xi=3$. The insert shows the probability density function (PDF) distribution of $f_{C}$ within the steady states. (b) $f_{C}$ as a function of the probability $\rho$ for $\xi=0,1,...,4$. The two insets show the zoom-in of the two transition regions. (c) Phase diagram for $f_{C}$ in the parameter domain $\rho-b$ with $\xi=3$. Other parameter: $b=0.1$ in (a,b).
@@ IV.2 Probabilistic mixing
Figure 8: Evolution of Q-tables and state densities in PM. Left column: The temporal evolution of averaged Q-value difference $\overline{\Delta Q}_{s_{j}}$ defined in Eq. (8) for all six states. Right column: The temporal evolution of all six state densities $f_{s_{0,...,5}}$. Parameter for the three rows: (a, d) $\rho=0.3$, (b, e) $\rho=0.4$, (c, f) $\rho=0.5$. Other parameter: $\xi=3$.
@@
$\bar{f}\left(s,s^{\prime}\right):=\frac{\sum_{\tau=t_{0}}^{t-1}\sum_{i=1}^{N}\mathbbm{1}_{s^{i}(\tau)=s,s^{i}(\tau+1)=s^{\prime}}}{N\left(t-t_{0}-1\right)}.$ (10)
Here the function $\mathbbm{1}_{\text{predicate}}$ equals one if the predicate of the two consecutive states for player $i$ is true, and zero otherwise. Therefore, Cohen’s kappa coefficient $\kappa(s,s^{\prime})$ captures the transition “flow” from state $s$ to $s^{\prime}$. If a pair of non-diagonal positive coefficients $\kappa(s,s^{\prime})\approx\kappa(s^{\prime},s)$, this corresponds to a detailed balance, otherwise there is a net transition flow between the two states. Note that diagonal elements of the matrix are for self-loops, where the system could settle down.
@@ VI Conclusion and discussion
In both mixing schemes, Q-learning players act as a “cooperation lubricant” in scenarios where $\xi=3$, where cooperation would otherwise fade due to the inability of hard conditional cooperators alone to sustain it. Remarkably, through Q-learning, players cooperate to achieve higher long-term rewards, ultimately fostering a significant level of cooperation in environments that would typically favor defection. Notice that researchers are recently aware of the behavioral multi-modality in humans, where individuals opt for different behavioral modes at various scenarios, and start to reveal new complexities therein Han et al. (2022); Ma et al. (2023); Chen and Fu (2023); Sheng et al. (2024).